# VGC-Bench: A Benchmark for Generalizing Across Diverse Team Strategies in Competitive Pokémon

### v2

@@ Appendix A Game Analysis
Pokémon VGC is an incredibly complex game; there are up to four Pokémon out at a time, each able to target up to three targets with a move and potentially terastallize or switch out into up to two benched Pokémon. Then we can calculate the approximate branching factor of nondeterminism in Pokémon moves. Calculated loosely, there are 16 damage rolls for each damaging move, almost always with a chance to hit normally, fail, land a critical hit, land an additional effect, or both land a critical hit and an additional effect. If we only consider these possibilities, there are outcomes for every move used. Also, each Pokémon has 4 total moves they can use, each of which have a maximum of 3 possible targets. If terastallization is available, there are
@@
Additionally, we can quantify the worst-case size of a player’s information set, or the number of indistinguishable states due to Pokémon VGC’s partial observability. The only unobservables for a player are the opponent Pokémon’s stats and the Pokémon that haven’t yet been revealed to be brought by the opponent during team preview. For the latter, the opponent immediately reveals half of their team preview decision at the beginning of the battle, leaving only two unrevealed Pokémon left out of 4 possibilities. That leaves possibilities. Pokémon stats are determined by a Pokémon’s EV’s, IV’s, and natures. There are 25 natures (although 5 are identical in effect, so we will use 21), 32 possible values for each of its 6 IV values, and 512 points to distribute over 6 stats for EV values, with no more than 252 points for each stat. Assuming all EV points are used and all EV spreads would distribute points to stats that are divisible by 4, which is reasonable because EV points only affect stats in multiples of 4, we can formalize EV spreading as the problem of distributing units across 6 stats, with the maximum allocation per stat being . We must find the total number of solutions to the equation , for . This can be solved using the inclusion-exclusion principle for bounded integer compositions, and yields a total count of 246,774,528. Additionally, players almost always set all of their Pokémon’s IVs to 31, so we will leave the IV configuration space out of the calculation so as to not overestimate the realistic size of the configuration space. Thus, the number of possible realistic stat configurations is
@@
We will now compare this with the initial configuration space of other games. In chess and Go, players can only initialize as white or black, and in Jeopardy there is no initial configuration at all. In Texas hold ’em Poker, each player is dealt 2 cards out of a 52-card deck at the beginning of the game, yielding a total initialization space of for the first player, for the second player, etc., or a total configuration space of
Dota has two teams of five players choose from a total of 126 available heroes, yielding a total configuration space of . Finally, Starcraft allows each of the two players to choose from three available races, yielding total configurations.
Appendix B ELO Ratings
To calculate ELO ratings from a matrix of win rates for , we define the implied rating difference via the formula
so that
We then form the linear system
for all pairs with valid comparisons, and add the constraint
to fix the mean rating. We solve this overdetermined system in the least-squares sense to obtain preliminary ratings . Finally, we shift all ratings so that
and, if desired, rescale so that the standard deviation satisfies
Appendix C Numerical Cross-Play Results
| Algorithm | R | MBP | SH | LLM | SP | FP | DO | BC | BCSP | BCFP | BCDO |
| R | – | 6 | 1 | 38 | 1 | 2 | 4 | 5 | 0 | 1 | 4 |
| MBP | 94 | – | 40 | 88 | 29 | 48 | 33 | 53 | 22 | 6 | 27 |
| SH | 99 | 60 | – | 91 | 52 | 40 | 100 | 65 | 25 | 22 | 62 |
| LLM | 62 | 12 | 9 | – | 5 | 2 | 2 | 12 | 1 | 0 | 05 |
| SP | 99 | 71 | 48 | 95 | – | 30 | 95 | 73 | 20 | 61 | 62 |
| FP | 98 | 52 | 60 | 98 | 70 | – | 38 | 88 | 63 | 83 | 78 |
| DO | 96 | 67 | 0 | 98 | 05 | 62 | – | 83 | 40 | 17 | 08 |
| BC | 95 | 47 | 35 | 88 | 27 | 12 | 17 | – | 13 | 04 | 18 |
| BCSP | 100 | 78 | 75 | 99 | 80 | 37 | 60 | 87 | – | 38 | 55 |
| BCFP | 99 | 94 | 78 | 100 | 39 | 17 | 83 | 96 | 62 | – | 89 |
| BCDO | 96 | 73 | 38 | 95 | 38 | 22 | 92 | 82 | 45 | 11 | – |
| Algorithm | R | MBP | SH | LLM | SP | FP | DO | BC | BCSP | BCFP | BCDO |
| R | – | 9 | 1 | 48 | 3 | 2 | 4 | 16 | 0 | 2 | 0 |
| MBP | 91 | – | 38 | 85 | 25 | 21 | 35 | 66 | 8 | 12 | 10 |
| SH | 99 | 62 | – | 97 | 35 | 45 | 53 | 79 | 34 | 32 | 21 |
| LLM | 52 | 15 | 3 | – | 4 | 2 | 7 | 16 | 3 | 2 | 0 |
| SP | 97 | 75 | 65 | 96 | – | 53 | 69 | 90 | 38 | 46 | 35 |
| FP | 98 | 79 | 55 | 98 | 47 | – | 53 | 87 | 17 | 38 | 33 |
| DO | 96 | 65 | 47 | 93 | 31 | 47 | – | 74 | 14 | 18 | 14 |
| BC | 84 | 34 | 21 | 84 | 10 | 13 | 26 | – | 5 | 7 | 10 |
| BCSP | 100 | 92 | 66 | 97 | 62 | 83 | 86 | 95 | – | 63 | 59 |
| BCFP | 98 | 88 | 68 | 98 | 54 | 62 | 82 | 93 | 37 | – | 58 |
| BCDO | 100 | 90 | 79 | 100 | 65 | 67 | 86 | 90 | 41 | 42 | – |
| Algorithm | R | MBP | SH | LLM | SP | FP | DO | BC | BCSP | BCFP | BCDO |
| R | – | 10 | 1 | 41 | 1 | 2 | 5 | 13 | 3 | 1 | 0 |
| MBP | 90 | – | 23 | 93 | 28 | 28 | 32 | 59 | 18 | 11 | 7 |
| SH | 99 | 77 | – | 91 | 48 | 61 | 48 | 73 | 33 | 24 | 25 |
| LLM | 59 | 7 | 9 | – | 1 | 7 | 3 | 14 | 1 | 2 | 2 |
| SP | 99 | 72 | 52 | 99 | – | 48 | 67 | 68 | 35 | 28 | 28 |
| FP | 98 | 72 | 39 | 93 | 52 | – | 64 | 84 | 37 | 29 | 21 |
| DO | 95 | 68 | 52 | 97 | 33 | 36 | – | 63 | 35 | 19 | 20 |
| BC | 87 | 41 | 27 | 86 | 32 | 16 | 37 | – | 10 | 12 | 13 |
| BCSP | 97 | 82 | 67 | 99 | 65 | 63 | 65 | 90 | – | 51 | 40 |
| BCFP | 99 | 89 | 76 | 98 | 72 | 71 | 81 | 88 | 49 | – | 52 |
| BCDO | 100 | 93 | 75 | 98 | 72 | 79 | 80 | 87 | 60 | 48 | – |
| Algorithm | R | MBP | SH | LLM | SP | FP | DO | BC | BCSP | BCFP | BCDO |
| R | – | 11 | 2 | 45 | 5 | 6 | 4 | 11 | 4 | 2 | 2 |
| MBP | 89 | – | 19 | 92 | 36 | 28 | 31 | 60 | 21 | 13 | 22 |
| SH | 98 | 81 | – | 94 | 49 | 57 | 42 | 80 | 33 | 38 | 26 |
| LLM | 55 | 8 | 6 | – | 7 | 1 | 6 | 16 | 5 | 3 | 2 |
| SP | 95 | 64 | 51 | 93 | – | 45 | 54 | 73 | 33 | 31 | 27 |
| FP | 94 | 72 | 43 | 99 | 55 | – | 44 | 67 | 32 | 33 | 39 |
| DO | 96 | 69 | 58 | 94 | 46 | 56 | – | 66 | 28 | 26 | 31 |
| BC | 89 | 40 | 20 | 84 | 27 | 33 | 34 | – | 15 | 10 | 17 |
| BCSP | 96 | 79 | 67 | 95 | 67 | 68 | 72 | 85 | – | 48 | 45 |
| BCFP | 98 | 87 | 62 | 97 | 69 | 67 | 74 | 90 | 52 | – | 49 |
| BCDO | 98 | 78 | 74 | 98 | 73 | 61 | 69 | 83 | 55 | 51 | – |
Appendix D Misc.
| Element Range | Action Description |

### v3

@@ Appendix A Game Analysis
Pokémon VGC is a complex game; there are up to four Pokémon out at a time, each able to target up to three targets with a move and potentially terastallize or switch out into up to two benched Pokémon. Then we can calculate the approximate branching factor of nondeterminism in Pokémon moves. Calculated loosely, there are 16 damage rolls for each damaging move, almost always with a chance to hit normally, fail, land a critical hit, land an additional effect, or both land a critical hit and an additional effect. If we only consider these possibilities, there are outcomes for every move used. Also, each Pokémon has 4 total moves they can use, each of which have a maximum of 3 possible targets. If terastallization is available, there are
@@
Additionally, we can quantify the worst-case size of a player’s information set, or the number of indistinguishable states due to Pokémon VGC’s partial observability. The only unobservables are the opponent Pokémon’s stats and the 2 bench Pokémon the opponent selected during team preview. For the latter, there are possibilities. As for stats, Pokémon stats are determined by a Pokémon’s EV’s, IV’s, and natures. There are 25 natures (although 5 are identical in effect, so we will use 21), 32 possible values for each of its 6 IV values, and 512 points to distribute over 6 stats for EV values, with no more than 252 points for each stat. Assuming all EV points are used and all EV spreads would distribute points to stats that are divisible by 4, which is reasonable because EV points only affect stats in multiples of 4, we can formalize EV spreading as the problem of distributing units across 6 stats, with the maximum allocation per stat being . We must find the total number of solutions to the equation , for . This can be solved using the inclusion-exclusion principle for bounded integer compositions, and yields a total count of 246,774,528. We will leave IVs out of the calculation since most Pokémon are given all 31 points per stat. Thus, the number of possible realistic stat configurations is
@@
In chess and Go, players can only initialize as white or black. In Texas hold ’em Poker, each player is dealt 2 cards out of a 52-card deck at the beginning of the game, yielding a total initialization space of . Dota has two teams of five players choose from a total of 126 available heroes, yielding a total configuration space of . Finally, Starcraft allows each of the two players to choose from three available races, as well as 9 playable maps, yielding total configurations.
Appendix B Additional Tables
| Index | Action Description |
@@ Appendix B Additional Tables
Algorithm R MBP SH LLM SP FP DO BC BCSP BCFP BCDO R – 0.068 0.039 0.220 0.007 0.014 0.008 0.053 0.001 0.002 0.012 MBP 0.932 – 0.298 0.660 0.134 0.070 0.116 0.440 0.064 0.050 0.136 SH 0.961 0.702 – 0.820 0.215 0.196 0.229 0.551 0.110 0.091 0.178 LLM 0.780 0.340 0.180 – 0.110 0.080 0.070 0.310 0.020 0.020 0.060 SP 0.993 0.866 0.785 0.890 – 0.587 0.707 0.790 0.229 0.432 0.448 FP 0.986 0.930 0.804 0.920 0.413 – 0.482 0.805 0.229 0.239 0.266 DO 0.992 0.884 0.771 0.930 0.293 0.518 – 0.842 0.203 0.236 0.337 BC 0.947 0.560 0.449 0.690 0.210 0.195 0.158 – 0.048 0.025 0.126 BCSP 0.999 0.936 0.890 0.980 0.771 0.771 0.797 0.952 – 0.510 0.744 BCFP 0.998 0.950 0.909 0.980 0.568 0.761 0.764 0.975 0.490 – 0.556 BCDO 0.988 0.864 0.822 0.940 0.552 0.734 0.663 0.874 0.256 0.444 –
Algorithm R MBP SH LLM SP FP DO BC BCSP BCFP BCDO R – 0.079 0.041 0.210 0.016 0.023 0.013 0.036 0.003 0.002 0.007 MBP 0.921 – 0.369 0.760 0.187 0.182 0.176 0.407 0.091 0.094 0.098 SH 0.959 0.631 – 0.750 0.289 0.262 0.302 0.518 0.132 0.143 0.133 LLM 0.790 0.240 0.250 – 0.140 0.060 0.080 0.290 0.060 0.030 0.020 SP 0.984 0.813 0.711 0.860 – 0.513 0.488 0.671 0.246 0.239 0.263 FP 0.977 0.818 0.738 0.940 0.487 – 0.569 0.647 0.273 0.293 0.253 DO 0.987 0.824 0.698 0.920 0.512 0.431 – 0.653 0.222 0.256 0.213 BC 0.964 0.593 0.482 0.710 0.329 0.353 0.347 – 0.130 0.114 0.171 BCSP 0.997 0.909 0.868 0.940 0.754 0.727 0.778 0.870 – 0.523 0.505 BCFP 0.998 0.906 0.857 0.970 0.761 0.707 0.744 0.886 0.477 – 0.522 BCDO 0.993 0.902 0.867 0.980 0.737 0.747 0.787 0.829 0.495 0.478 –
Algorithm R MBP SH LLM SP FP DO BC BCSP BCFP BCDO R – 0.067 0.052 0.270 0.040 0.037 0.032 0.038 0.010 0.003 0.005 MBP 0.933 – 0.337 0.680 0.310 0.272 0.274 0.361 0.104 0.089 0.084 SH 0.948 0.663 – 0.770 0.413 0.358 0.432 0.478 0.152 0.160 0.147 LLM 0.730 0.320 0.230 – 0.200 0.110 0.140 0.210 0.070 0.080 0.020 SP 0.960 0.690 0.587 0.800 – 0.479 0.474 0.551 0.209 0.208 0.241 FP 0.963 0.728 0.642 0.890 0.521 – 0.486 0.578 0.211 0.232 0.216 DO 0.968 0.726 0.568 0.860 0.526 0.514 – 0.533 0.228 0.215 0.200 BC 0.962 0.639 0.522 0.790 0.449 0.422 0.467 – 0.204 0.191 0.193 BCSP 0.990 0.896 0.848 0.930 0.791 0.789 0.772 0.796 – 0.498 0.481 BCFP 0.997 0.911 0.840 0.920 0.792 0.768 0.785 0.809 0.502 – 0.486 BCDO 0.995 0.916 0.853 0.980 0.759 0.784 0.800 0.807 0.519 0.514 –
Algorithm R MBP SH LLM SP FP DO BC BCSP BCFP BCDO R – 0.081 0.029 0.210 0.054 0.053 0.066 0.050 0.008 0.010 0.006 MBP 0.919 – 0.344 0.670 0.386 0.388 0.407 0.398 0.122 0.127 0.117 SH 0.971 0.656 – 0.790 0.486 0.482 0.490 0.511 0.166 0.173 0.199 LLM 0.790 0.330 0.210 – 0.240 0.120 0.190 0.150 0.060 0.040 0.030 SP 0.946 0.614 0.514 0.760 – 0.512 0.510 0.481 0.169 0.189 0.196 FP 0.947 0.612 0.518 0.880 0.488 – 0.487 0.466 0.198 0.180 0.160 DO 0.934 0.593 0.510 0.810 0.490 0.513 – 0.508 0.197 0.189 0.198 BC 0.950 0.602 0.489 0.850 0.519 0.534 0.492 – 0.231 0.234 0.238 BCSP 0.992 0.878 0.834 0.940 0.831 0.802 0.803 0.769 – 0.554 0.505 BCFP 0.990 0.873 0.827 0.960 0.811 0.820 0.811 0.766 0.446 – 0.469 BCDO 0.994 0.883 0.801 0.970 0.804 0.840 0.802 0.762 0.495 0.531 –